OpenStudy (anonymous):

Cards with the numbers 1 through 50 are placed in a box and thoroughly shuffled. What is the probability of randomly selecting a card with a number less than 10 or a multiple of 12?

OpenStudy (anonymous):

count

OpenStudy (anonymous):

there are 9 numbers less than 10 how many others are multiples of 12?

OpenStudy (anonymous):

\[\frac{ 9 }{ 50 }\]and\[\frac{ 4 }{ 50}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\frac{ 13 }{ 50 }\]

OpenStudy (anonymous):

oh i see what you are saying 12,24,36,48 makes 4 for a total of \(13\) yes, that looks good to me \(\frac{13}{50}\)

OpenStudy (anonymous):

Thank you ^^

OpenStudy (anonymous):

Refer to the attachment.